there are at most eight marbles in a bag, you pick out two red marbles and are told the probability of picking out two red marbles is 50%. How many marbles are there and exactly how many marbles are red?
Answer #1

let there are x red marbles and total y are total marbles

here x<y ;x>=2 and y<=8

therefore probability of picking out two red marbles =(x/y)*(x-1)/(y-1) =0.5

x2-x =0.5*(y-y2)

checking above for values of 3<=y<=8

total number of marbles =4

and total number of red marbles =3